OpenStudy (awesome781):

A softball with a mass of 0.11 kg moves at a speed of 12 m/s. Then the ball is hit by a bat and rebounds in the opposite direction at a speed of 15 m/s . What is the change in momentum of the ball?

OpenStudy (wach):

Do you know what momentum p is equal to?

OpenStudy (wach):

Also I assume 'bullet' is a ball actually?

OpenStudy (awesome781):

It is softball.... sorry typo

OpenStudy (awesome781):

No I don't know what momentum p is.....

OpenStudy (wach):

momentum = p = mv = mass * velocity Mass is constant but the force of the bat causes the ball to accelerate to a different velocity When finding change in something, we find the difference or change in magnitude and divide that by the initial magnitude

OpenStudy (awesome781):

Ok

OpenStudy (wach):

The initial momentum = m1v1 = 0.11kg*12m/s = 1.32 Final = m1v2 = 0.11*15 = 1.65 Change = 1.65 - 1.32 = 0.33 0.33 kg*m/s is the change in momentum

OpenStudy (awesome781):

Thank you

OpenStudy (wach):

If you need it as a percentage divide 0.33 by 1.65 no problem :)

OpenStudy (awesome781):

Ok...I got it!